How many square meters of fabric are needed to make 120 triangular pennants each with a base of 30 cm and a height of 48 cm
Found 2 solutions by macston, Alan3354:
Answer by macston(5194)

.
b=base=30cm=0.30m; h=height=48cm=0.48m
area of each pennant=(1/2)bh
Total area needed=(120)(1/2)(0.30m)(0.48m)=8.64 square meters

Answer by Alan3354(69443)

How many square meters of fabric are needed to make 120 triangular pennants each with a base of 30 cm and a height of 48 cm
----------------
The area of each triangle = b*h/2 = 30*48/2 = 720 sq cm
---------------------
1 sq meter = 10,000 sq cms